OpenStudy (sidsiddhartha):

the characteristic equation of a (3X3) matrix P is defined as \[a(\lambda)=|\lambda I-P|=\lambda^3+\lambda^2+2 \lambda+I=0\] if I denotes identity matrix then find the inverse of P

OpenStudy (ikram002p):

*

OpenStudy (sidsiddhartha):

;}

OpenStudy (sidsiddhartha):

im thinking about carey hamilton

OpenStudy (sidsiddhartha):

every square matrix satisfies its characteristic equation so can i substitute \[\lambda=P\]

ganeshie8 (ganeshie8):

\[\large P^3+P^2+2P+I = 0\] like this ?

OpenStudy (sidsiddhartha):

yes it can be written right?

OpenStudy (sidsiddhartha):

@ganeshie8

ganeshie8 (ganeshie8):

definitely thats what cayley-hamilton theorem is about, right ?

OpenStudy (sidsiddhartha):

yeah now if i multiply P^-1 in both sides i think i'll get a result

ganeshie8 (ganeshie8):

looks good to me but i need to look my notes lol, not really sure on these..

OpenStudy (sidsiddhartha):

yeah mee too i've done this 3 years ago and almost forgot everything lol

ganeshie8 (ganeshie8):

\[\large P^3+P^2+2P = -I\]

ganeshie8 (ganeshie8):

\[\large P^{-1}(P^3+P^2+2P = -I)\]

ganeshie8 (ganeshie8):

looks fine to me :)

OpenStudy (sidsiddhartha):

yeah \[P^{-1}=-(p^2+P+2I)\]

OpenStudy (sidsiddhartha):

yeah its one of the option okay thanks :)

ganeshie8 (ganeshie8):

nice :D